OpenStudy (mckenzieandjesus):

Find the common ratio of the sequence. –4, 8, –16, 32, . . . –2 –12 -1/-2 12

Nnesha (nnesha):

to find common ratio divide next term by the previous one \[\large\rm r=\frac{ a_2 }{ a_1}~~or~~\frac{ a_3 }{ a_2} \]a_2 = 2nd term a_3 = 3rd term

OpenStudy (mckenzieandjesus):

wouldnt the next term be 36?

OpenStudy (mckenzieandjesus):

i'm getting -9 but thats not right..

Nnesha (nnesha):

ratio between two numbers in geometric you will multiply the first by common ratio to get the nexxt one other way to solve for r is first term times r = 2nd term

Nnesha (nnesha):

\[\huge\rm a_1 \times r = a_2\] what is first term and what is 2nd one ?

Nnesha (nnesha):

aww i see

OpenStudy (mckenzieandjesus):

-4 and 36

Nnesha (nnesha):

i see what u did there it's `common ratio` not common difference

Nnesha (nnesha):

now first one is -4 2nd term is what ?

Nnesha (nnesha):

1st term , 2nd term , 3rd, 4th term..........

OpenStudy (mckenzieandjesus):

32?

Nnesha (nnesha):

2 comes after 1 so what's 2nd term in this sequence `–4, 8, –16, 32, . . .`

OpenStudy (mckenzieandjesus):

8

Nnesha (nnesha):

right so \[\huge\rm -4r=8\] solve for r

OpenStudy (mckenzieandjesus):

2?

Nnesha (nnesha):

hmm positive 8 divide by `negative 4` = ?

OpenStudy (mckenzieandjesus):

-32

Nnesha (nnesha):

divide not multiply

OpenStudy (mckenzieandjesus):

ohh

OpenStudy (mckenzieandjesus):

-2!

Nnesha (nnesha):

to find common ratio we should divide 2nd term by first one or 3rd term divide by 2nd term

Nnesha (nnesha):

right yahoo!

OpenStudy (mckenzieandjesus):

so i gueesed the right answer then :) thanku

OpenStudy (mckenzieandjesus):

can u help me with another?

OpenStudy (mckenzieandjesus):

2, –10, 50, –250, . . .

Nnesha (nnesha):

will try

Nnesha (nnesha):

find common ratio ?

OpenStudy (mckenzieandjesus):

yes

Nnesha (nnesha):

alright same steps try it!

Nnesha (nnesha):

divide any term with the previous one you can divide 4th by 3rd one 3rd one by 2nd one 2nd one by first one